OpenStudy (anonymous):

Solve for x in the following proportion: 3/5 = 18/x A. x = 10 B. x = 20 C. x = 30 D. x = 40

OpenStudy (anonymous):

3/5 = 18/x (cross multiply) (3)(x) = (18)(5) 3x = 90 x = 30